Peter starts with $10,000 and Paul with $1,000. They bet $100 each on each game. In each game, each player has the same chance of winning. What is the probability that Peter goes broke?
Answer #1

since probability is same for both the players and initial money for perter a =10000 and paul =1000

probability that Peter goes broke =1-a/(a+b) =1-10000/(10000+1000)= 0.0909